OpenStudy (anonymous):

double integral sec(x^2 +y^2) dx dy . Use polar coordinates

OpenStudy (anonymous):

Also, x^2+y^2=1

OpenStudy (anonymous):

0 to 1 ^

OpenStudy (blockcolder):

\[\iint_D \sec(x^2+y^2)\ dx\ dy\] What's the region of integration?

OpenStudy (anonymous):

x^2 +y^2=<1

OpenStudy (blockcolder):

Okay, so in polar coordinates, that region can be written like this: \[D=\{(r,\theta)|0\leq r\leq1, 0\leq \theta \leq 2\pi\}\] Then rewrite the double integral like this: \[\int_0^{2\pi} \int_0^1 \sec(r^2)\ r\ dr\ d\theta\] I know you can evaluate that. :)

OpenStudy (anonymous):

thanx. This one is tedious